# GraphMoE: Amplifying Cognitive Depth of Mixture-of-Experts Network via Introducing Self-Rethinking Mechanism

### v2

@@ 5.1 Evaluating the Performance of GraphMoE
The degree of enhancements varies across models, influenced by their specific base architectures. This variability is attributed to the integration level, or the combination mode, between the MoE and LoRA components, which can impact the parameter space for incremental learning and thereby modify the marginal benefits of GraphMoE. Specifically, as highlighted by Li et al. (2024), the MoLA and LoRAMoE configurations incorporate LoRA modules as expert models by integrating them with outputs from either the attention layers or feedforward network (FFN) layers—an approach termed the “Plugged-in” method. In contrast, MixLoRA employs a “fused” integration strategy, embedding the LoRA expert models directly within the original network layers. This design allows the LoRA experts to exert a more immediate and substantial influence on the conventional LLM components, resulting in the most significant enhancement of the original LoRA+MoE architecture.
@@ 5.2 Main Results
Table 4: Performance evaluation of LoRA-Based method SFT across various benchmarks. Results marked with an asterisk (*) are experimental data sourced from Li et al. (2024), conducted with full precision (FP32). All other results were obtained using BF16 precision. In the table, derivatives of baselines that employ the DoRA technique are indicated by replacing “L” with “D” in the original model name.
The experimental results are detailed in Table 4. Due to limitations in computational resources, the experiments conducted in this study employ reduced precision BF16 for both training and inference phases. However, we include the full precision results of MixLoRA as a reference for comparison. It is evident that all reproduced models experience a significant performance decline compared to their full precision counterparts. Nonetheless, our GraphMoE models with MixLoRA and MixDoRA as base models still achieve SOTA performance even compared to the full precision results. This substantial improvements compared to other SOTA LoRA+MoE models further demonstrate the effectiveness of the GraphMoE architecture.
@@ 5.4 Sensitivity Analysis and Overhead
Figure 5: Sensitivity analysis of the additional hyperparameters in the GraphMoE architecture. The impact on GraphMoE’s computational overhead is demonstrated by examining how the inference time scales with the Reasoning Round ($T$).

### v3

@@ 5.1 Evaluating the Performance of GraphMoE
The degree of enhancements varies across models, influenced by their specific base architectures. This variability is attributed to the integration level, or the combination mode, between the MoE and LoRA components, which can impact the parameter space for incremental learning and thereby modify the marginal benefits of GraphMoE. Specifically, as highlighted by Li et al. (2024), the MoLA and LoRAMoE configurations incorporate LoRA modules as expert models by integrating them with outputs from either the attention layers or MLP layers—an approach termed the “Plugged-in” method. In contrast, MixLoRA employs a “fused” integration strategy, embedding the LoRA expert models directly within the original network layers. This design allows the LoRA experts to exert a more immediate and substantial influence on the conventional LLM components, resulting in the most significant enhancement of the original LoRA+MoE architecture.
@@ 5.2 Main Results
Base (ICL-$3$) 0% 90.6 76.7 45.9 73.8 61.6 71.5 64.1 59.2 65.6
Base (ICL-$3$, Self-cons-55) 0% 91.1 79.4 50.2 73.7 57.0 70.3 66.9 62.1 66.3
Base (ICL-$3$, Self-cons-$10$) 0% 91.6 78.1 42.8 75.3 59.8 71.3 69.9 61.4 66.2
Base (ICL-$3$, Self-cons-$15$) 0% 91.9 78.0 45.6 72.0 67.9 71.7 63.3 63.2 66.4
@@
Table 4: Performance evaluation of LoRA-Based method SFT across various benchmarks. The term “ICL-n” indicates that n examples are included in the prompt for in-context learning, while “Self-cons-5” means that 5 samples are collected for self-consistency. “Param” refers to the size of the trainable parameters used in the experiments. Results marked with an asterisk (*) are experimental data sourced from Li et al. (2024), conducted with full precision (FP32). All other results were obtained using BF16 precision. In the table, derivatives of baselines that employ the DoRA technique are indicated by replacing “L” with “D” in the original model name.
The experimental results are detailed in Table 4. The initial part of the experiment involves four base language models (LLaMA-3-8B) as reference models that incorporate in-context learning and self-consistency techniques. In-context learning (ICL-n) refers to directly providing the task’s instructions in the prompt along with a certain number of demonstration examples. Self-consistency (Self-cons-n) refers to the common technique in LLMs where multiple reasoning paths are sampled and the most consistent answer is selected. Increasing the samples for self-consistency to 5, 10, and 15 yields performance scores of 66.3, 66.2, and 66.4, respectively, indicating minimal variation. Although self-consistency results in slight performance improvements beyond 65.6 on downstream tasks, the base language model still falls short when compared to those employing the LoRA-based SFT method.
Due to limitations in computational resources, the experiments conducted in this study employ reduced precision BF16 for both training and inference phases. However, we include the full precision results of MixLoRA as a reference for comparison. It is evident that all reproduced models experience a significant performance decline compared to their full precision counterparts. Nonetheless, our GraphMoE models with MixLoRA and MixDoRA as base models still achieve SOTA performance even compared to the full precision results. This substantial improvements compared to other SOTA LoRA+MoE models further demonstrate the effectiveness of the GraphMoE architecture.
@@ 5.4 Sensitivity Analysis and Overhead
Figure 5: Sensitivity analysis of the additional hyperparameters in the GraphMoE architecture. The impact on GraphMoE’s computational overhead is demonstrated by examining how the inference time scales with the Reasoning Round ($T$). In subfigure (c), “Infer. Time” refers to the inference time, with the duration of the first round set as a unit. The figure displays the factors by which the time of each subsequent round compares to that of the first round.
@@
5.5 Effect of Parameter Size
on Model Performance
Method Param ARC-E ARC-C BoolQ OBQA PIQA SIQA HellaS WinoG AVG
LoRA 1.9% 87.5 78.9 69.8 84.4 84.4 77.1 88.4 49.1 77.5
LoRA 2.6% 86.2 76.7 70.2 78.2 82.9 74.2 80.1 50.7 74.9
LoRA 5.9% 84.7 78.6 63.6 82.2 82.1 32.9 84.1 50.4 69.8
LoRA 7.9% 83.9 73.1 62.2 84.4 82.5 74.6 92.7 49.5 75.4
LoRA 9.9% 85.9 73.3 70.0 82.0 85.7 32.1 89.1 49.6 71.0
MixLoRA 3.0% 86.9 77.0 74.0 84.4 86.0 75.5 93.7 83.3 82.6
MixLoRA 6.0% 87.3 79.0 74.9 88.4 88.1 76.7 93.6 82.5 83.8
MixLoRA 9.0% 86.7 77.0 71.7 86.2 80.5 74.0 89.5 80.4 80.8
GraphMoE 5.9% 89.2 80.9 75.8 89.6 87.8 77.0 95.4 83.2 84.9
Table 5: Performance Comparison of Parameter Variants. In the table, GraphMoE refers to the GraphMoE(MixLoRA) model.
It has been claimed in previous works Tang et al. (2023b); Li et al. (2024) that simply increasing parameter size will not give a metric increase when doing SFT. We conduct an experiment to further demonstrate this phenomenon. From Table 5, it can be observed that increasing the parameter size does not consistently lead to improved performance. For instance, the LoRA method with 1.9% parameters achieves the best average score of 77.5, while increasing the parameter size to 7.9% results in an series of fluctuation below the score of 77.5. The similar phenomenon has been also observed in (b) of Figure 5, where the performance is not increasing when the GRU hidden size scaling factor increases.
Moreover, we directly increase one of the baseline model of GraphMoE to show if the increased performance was achieved by the additionally increased parameter size. Even though the performance of MixLoRA with 6.0% parameters results in a higher average score of 83.9, the larger size of Mixlora with 9.0% parameters result in an inferior performance. This indicates that the self-rethinking mechanism plays the pivotal role in effective parameter utilization and model structuring, as showcased by the comparison between MixLoRA and GraphMoE(MixLoRA), and implementing the GraphMoE framework can yield better results without proportionally increasing parameter size.
Overall, it can be observed that the best overall performance is observed with GraphMoE(MixLoRA), which uses just 5.9% of parameters yet achieves the highest average score of 84.9. This demonstrates that optimal model design principles, rather than mere parameter scaling, are crucial for achieving superior results in these tasks.
Method Param ARC-E ARC-C BoolQ OBQA PIQA SIQA HellaS WinoG AVG
[…]
MixLoRA 6.0% 87.3 79.0 74.9 88.4 88.1 76.7 93.6 82.5 83.8
$\Delta$ - -0.1 -0.2 1 3.2 2.8 3.8 -1.4 2.1 1.4
[…]
GraphMoE 5.9% 89.2 80.9 75.8 89.6 87.8 77.0 95.4 83.2 84.9
$\Delta$ - 1.1 1.9 1.7 1.8 2.2 2.7 2.2 -1.9 1.5
[…]
To provide further insights into the utilization of the MoE structure, we experimented with activating additional experts within the MoE layers to assess if similar performance levels could be achieved, and the results are presented in Table 6. Surprisingly, increasing the number of active experts led to a decline in performance. This finding underscores that additional parameters do not necessarily guarantee improved reasoning performance. Instead, the effectiveness of the designed mechanisms within the neural architecture plays a more critical role, as demonstrated by previous results in Table 5.
5.6 In-depth Analysis of Computation Cost
To provide a comprehensive analysis of the increase in time consumption during each rethinking round, we conducted a token-level inference time test for each module executed in the forward step. The results are presented in Figure 6. It is evident that the additional GRU module introduced to the Router contributes marginal running time, accounting for less than $5\%$. This indicates that the majority of increased latency for each additional Reasoning Round ( $T$ ) is primarily due to the additional inference step of MLP module. The time ratios of the Attention Module and the MLP Module determine the slope of the lines shown in (c) of Figure 5. Two key points are highlighted to understand the relationship between module time latency and the inference time latency for each Reasoning Round ( $T$ ):
  • As the input sequence length increases, the time increase for the Attention Module occurs more rapidly than for the MLP Module. This is because the time complexity of the attention mechanism is $O(N^{2})$ [^8] NN denotes the length of the input sequence. due to the multiplication of the Query matrix $Q$ and the Key matrix KK [^9] For a detailed description, please refer to Vaswani et al. (2017).. The time complexity of the MLP, being a linear layer, is ( $O(N)$ ). However, due to the implementation of the KV Cache, which includes locally stored cache for historical attention key and value tensors of each layer, the actual time complexity of the Attention Module does not increase as mush as double. For a given task, a longer input sequence length results in a relatively smaller increase in the time for Reasoning Round ( $T$ ).
  • The different architectures of the MoE LoRA base models influence the time ratio observed in Figure 6. As detailed in Li et al. (2024), LoRAMoE features vanilla attention LoRA layers along with plugged MLP LoRA layers; MoLA includes plugged attention LoRA layers and plugged MLP LoRA-MoE layers; and MixLora comprises fused attention and fused MLP LoRA-MoE layers. Consequently, during inference, the time spent on MLP LoRA layers is highest for LoRAMoE, followed by MixLoRA, and then MoLA.
Figure 6: Comparison of the latency compositiong ratio of different based on different LoRA+MoE and GraphMoE architectures.
@@ 6 Conclusion
Appendix A Discussions
  • What is the value of $k$ in the top-$k$ selection? How is it selected? Currently, $k$ is treated as a hyperparameter, with two main considerations. First, aligning with other MoE architectures, the number of selected experts typically maintains fixed and sparse, with 2 out of 8 being a common configuration. Second, our framework facilitates expert collaboration, so the top $k$ is not a critical parameter for fully leveraging expert potential.
  • Is the top-k selection in the routing operation (Equation 4) not differentiable? Regarding the differentiability issue of the top-$k$ in Equation 4, it is important to note that while the top-$k$ operation itself is non-differentiable, it does not directly impact the gradient calculations. Rather, the fixed number of experts chosen through the top-k operation contribute to the subsequent gradient computations.
  • Are some experts learning to do the initial processing, while others more frequently do the later processing? In this study, the MoE is applied within each transformer block, so the experts do not possess specific physical significance. Therefore, we have not conducted an analysis of individual expert usage patterns or preferences within the reasoning process.